# SCREP: Scene Coordinate Regression and Evidential Learning-based Perception-Aware Trajectory Generation

### v1

@@ IV-B Trajectory Optimization
After receiving the current UAV pose from a fixed-lag smoother, reference waypoints, scene coordinates and their uncertainty, position and yaw trajectories are optimized to track waypoints and orient reliable scene coordinates. In this subsection, we detail the perception‑aware trajectory optimization.
@@
where the knot vector is $\{t_{0},t_{1},\dots,t_{n+k+1}\}$, the control points are $\mathbf{q}_{i}=\{q_{x,i},q_{y,i},q_{z,i},q_{\psi,i}\}\in\mathbb{R}^{4}$, and $N_{i,k}(t)$ denotes the B‑spline basis functions. Each knot span has equal duration $\Delta t=t_{i+1}-t_{i}$. For clamping, the first and last $k$ knots are repeated: $t_{0}=\dots=t_{k-1}$ and $t_{n+1}=\dots=t_{n+k+1}$. We fix the first control point to the current UAV pose and the last control point to the final waypoint within the planning horizon $T_{plan}$. Then, the other control points become optimization variables to minimize the cost function.
Optimization Cost. Similar to Bartolomei et al. [10], the total cost function, $C_{total}$, is represented as the weighted sum of the cost functions for different purposes. To reduce computational load, we optimize the position and yaw separately by minimizing two disjoint sub-objectives rather than the full $C_{total}$:
$\begin{gathered}C_{total}^{\mathbf{p}}=\lambda_{wp}C_{wp}^{\mathbf{p}}+\lambda_{eq}C_{eq}^{\mathbf{p}}+\lambda_{ie}C_{ie}^{\mathbf{p}}+\lambda_{s}C_{s}^{\mathbf{p}}\\ C_{total}^{\psi}=\lambda_{fov}C_{fov}^{\psi}+\lambda_{eq}C_{eq}^{\psi}+\lambda_{ie}C_{ie}^{\psi}+\lambda_{s}C_{s}^{\psi}.\end{gathered}$ (5)
where the superscripts, $\mathbf{p}$ and $\psi$, indicate that the terms are evaluated on the position and yaw trajectory, respectively. We are going to drop the superscript for brevity. The coefficients $\lambda_{(\cdot)}$ balance five components: $C_{wp}$ penalizes deviation from the given waypoints, $C_{fov}$ rewards keeping reliable scene coordinates within the camera’s field-of-view (FOV), $C_{eq}$ enforces the trajectory starts from the given initial velocity, acceleration, and jerk, $C_{ie}$ enforces the specified velocity and acceleration bounds, and $C_{s}$ promotes smoothness of the trajectory.
Let $\{\mathbf{w}_{0},\dots,\mathbf{w}_{r-1}\}$ denote all waypoints and let $\{s_{1},\dots,s_{r-2}\}$ be their prescribed arrival times except the first and last. The waypoint cost $C_{wp}$ penalizes the squared Euclidean error between the position trajectory and each desired waypoint:
$C_{wp}=\sum_{i=1}^{r-2}\,\sum_{o\in\{x,y,z\}}\bigl{(}q_{o}(s_{i})-w_{o,i}\bigr{)}^{2}.$ (6)
Our formulation of the FOV cost $C_{fov}$ is inspired by Murali et al. [3]. Whereas the original cost encourages the camera to track as many covisible features as possible within its frustum, our version instead rewards keeping low‑uncertainty scene coordinates in view regardless of covisibility.
@@
Next, we need to build a differentiable indicator function of whether scene coordinates $\mathbf{v}^{C}_{i,j}$ are within the camera frustum for the numerical optimization. For this purpose, the displacement vectors from the optical center to the four image‑plane corners:
$\begin{split}\mathbf{z}_{TR}&=\big{[}(w-c_{x})l_{x}\quad(-c_{y})l_{y}\quad f\big{]}^{T}\\ \mathbf{z}_{LR}&=\big{[}(w-c_{x})l_{x}\quad(h-c_{y})l_{y}\quad f\big{]}^{T}\\ \mathbf{z}_{TL}&=\big{[}(-c_{x})l_{x}\quad(-c_{y})l_{y}\quad f\big{]}^{T}\\ \mathbf{z}_{LL}&=\big{[}(-c_{x})l_{x}\quad(h-c_{y})l_{y}\quad f\big{]}^{T}\end{split}$ (7)
where $w$ and $h$ are the image width and height in pixels, $(c_{x},c_{y})$ is the principal point in pixels, $l_{x}$ and $l_{y}$ are the physical pixel sizes $(m/\text{pix})$, and $f$ is the focal length in meters. The normal vectors of the four side half planes of the camera frustum are $\mathbf{z}_{\mathrm{TR}}\!\times\!\mathbf{z}_{\mathrm{LR}}$, $\mathbf{z}_{\mathrm{TL}}\!\times\!\mathbf{z}_{\mathrm{TR}}$, $\mathbf{z}_{\mathrm{LL}}\!\times\!\mathbf{z}_{\mathrm{TL}}$, and $\mathbf{z}_{\mathrm{LR}}\!\times\!\mathbf{z}_{\mathrm{LL}}$. Using these normals, we construct a soft, indicator vector
$\mathbf{o}_{i,j}=\begin{bmatrix}\frac{1}{2}\!\bigl{[}1+\tanh\!\bigl{(}(\mathbf{z}_{\mathrm{TR}}\!\times\!\mathbf{z}_{\mathrm{LR}})\!\cdot\!\mathbf{v}^{C}_{i,j}/s\bigr{)}\bigr{]}\\[2.0pt] \frac{1}{2}\!\bigl{[}1+\tanh\!\bigl{(}(\mathbf{z}_{\mathrm{TL}}\!\times\!\mathbf{z}_{\mathrm{TR}})\!\cdot\!\mathbf{v}^{C}_{i,j}/s\bigr{)}\bigr{]}\\[2.0pt] \frac{1}{2}\!\bigl{[}1+\tanh\!\bigl{(}(\mathbf{z}_{\mathrm{LL}}\!\times\!\mathbf{z}_{\mathrm{TL}})\!\cdot\!\mathbf{v}^{C}_{i,j}/s\bigr{)}\bigr{]}\\[2.0pt] \frac{1}{2}\!\bigl{[}1+\tanh\!\bigl{(}(\mathbf{z}_{\mathrm{LR}}\!\times\!\mathbf{z}_{\mathrm{LL}})\!\cdot\!\mathbf{v}^{C}_{i,j}/s\bigr{)}\bigr{]}\\[2.0pt] \frac{1}{2}\!\bigl{[}1+\tanh\!\bigl{(}(\mathbf{v}^{C}_{i,j}-f\mathbf{e}_{3})\!\cdot\!\mathbf{e}_{3}/s\bigr{)}\bigr{]}\end{bmatrix},$ (8)
where $\mathbf{e}_{3}=[0,0,1]^{\mathsf{T}}$ and $s>0$ are the camera’s forward axis and a smoothing constant, respectively. For the $j$-th scene coordinate at time $t=s_{i}$, the desired scalar indicator is defined as the product of each element of $\mathbf{o}_{i,j}$ as
@@
$C_{fov}=-\sum_{i=1}^{r-1}\sum_{j=1}^{n_{f}}\exp(-aH_{\theta,j})F(\mathbf{c}_{i},\mathbf{v}_{j}).$ (10)
Here, $n_{f}$ denotes the number of selected scene coordinates. These coordinates are randomly chosen among those with entropy values below a predefined threshold. The weight $\exp(-aH_{\theta,j})$ depends on the entropy $H_{\theta,j}$ on Eq. (3); $a>0$ is constant. Lower entropy, i.e., higher certainty, yields a larger weight, encouraging the yaw trajectory to keep such scene coordinates in view.
@@
where $v_{o}(s_{i})$ and $a_{o}(s_{i})$ denote, respectively, the velocity and acceleration of coordinate $o$ at sample time $s_{i}$. The single auxiliary function, $M_{x}(x)=\text{ReLU}\big{(}(x^{2}-x_{\max}^{2})^{2}\big{)}$, is applied to both velocity and acceleration.
@@
TABLE I: Localization accuracy for each trajectory (translation / rotation in cm/). Columns group the three yaw strategies, Constant, Forward, and our perception‑aware planner (Ours), and distinguish when the accuracy is evaluated solely by SCR or the fixed‑lag smoother. Failed cases are not recorded and therefore are not included in the average calculation.
@@ V Experiments
We evaluated our approach under simulation with two goals: (a) to verify that the entropy defined in Eq. (3) provides an appropriate measure of uncertainty, and (b) to demonstrate the effectiveness of our perception‑aware trajectory planner. First, we empirically examined how the proposed entropy compares with conventional aleatoric and epistemic uncertainty measures. We then benchmarked our planner against baselines and performed an ablation study to isolate the impact of the uncertainty within SCR and the optimization. We also conducted the hardware-in-the-loop experiments to confirm the practical feasibility of our methods.
TABLE II: Parameters for trajectory optimization.
$\lambda_{wp}$ $\lambda_{fov}$ $\lambda_{eq}$ $\lambda_{ie}$ $\lambda_{s}$ $T_{\text{plan}}$ $T_{\text{exec}}$ $n_{f}$ $s$ $a$ $10^{4}$ $10^{1}$ $10^{3}$ 1.0 5.0 0.8s 0.5s 200 5.0 0.5
@@ V-A Setup
We ran photorealistic simulations using the FlightGoggles simulator [36] under Intel Core i9-7920X CPU. The simulated camera resolution was fixed at $480\times 480$ pixels, and its FoV was $90^{\circ}$. We trained a E‑SCRNet following the architecture in [6] for each of the two sites (“stata1” and “stata2”) in the stata_ground_floor environment as shown in Fig. 1.
To highlight the advantages brought by our perception‑aware planner, we deliberately used a small training set: about 1,200 images captured along a single trajectory (“picasso”) from the Blackbirds dataset [37], trained for only 100k iterations. The regularization coefficient $\rho$, the learning rates, and the batch size were $10^{-2}$, $10^{-4}$, and 1, respectively. Adam optimizer was applied. The training time was 14 minutes using one NVIDIA RTX 6000 Ada GPU.
For every planning step, we minimize the costs in Eq. (5) with the L-BFGS algorithm provided by NLopt [38]. The required parameters for the optimization are set in Table II. The number of control points for the clamped B-spline in Eq. (4) was set to six. Scene‑coordinate regression is run every $0.2\,\mathrm{s}$.
(a) L2 Error
(b) Density Histogram
Figure 4: Graphs for normalized uncertainty and scene coordinate errors of E-SCRNet. (a) Mean and standard deviation of L2 error (b) Density histogram
(a) Aleatoric Unc.
(b) Epistemic Unc.
(c) Entropy
(d) L2 Error
(e) Prediction
(f) Image
Figure 5: Qualitative results of E-SCRNet for a test image. (a) Aleatoric Uncertainty (b) Epistemic Uncertainty (c) Entropy (d) L2 Error of predicted scene coordinates (e) Predicted scene coordinates (f) Input image.
(a) Estimated trajectories & camera orientations
(b) Translation and rotation error
Figure 6: Comparison of the baseline planners (Constant, Forward) with the proposed method. (a) Top‑down view of the reference trajectory (black dotted), the executed trajectories (colored), camera orientation (green arrows). (b) Time histories of translation error (left) and rotation error (right) for Constant (blue), Forward (orange), and Ours (green).
@@ V-B Uncertainty Representation
We collected scene coordinate predictions and their uncertainty maps from E‑SCRNet while the UAV followed a circular trajectory. Fig. 4 analyzes the relationship between the L2 prediction error and three uncertainty estimates: aleatoric, epistemic, and entropy. We clipped the largest 3% of uncertainty values and normalized the remainder to $[0,1]$ for visual clarity.
As shown in Fig. 4(a), the mean error increases almost linearly with binned entropy, whereas its correlation with the others is weaker. Moreover, the error’s standard deviation declines as entropy decreases, while it remains high across the full range of the other measures. Fig. 4(b) further reveals that entropy is less skewed toward zero than the other uncertainties, indicating that it provides a more informative and well‑spread metric.
@@
Comparison to baselines. To the best of our knowledge, no prior work has explored perception‑aware trajectory generation using SCR. Therefore, we compare our planner with two intuitive yaw strategies: Constant, which fixes the camera to a global heading, and Forward, which keeps the camera aligned with the velocity vector. We evaluate four trajectories from the Blackbird dataset per site, “clover,” “patrick,” “thrice,” and “winter” in stata1, and “ampersand,” “dice,” “halfmoon,” and “sid” in stata2. Each trajectory is flown multiple times.
Table I reports localization accuracy for both the raw SCR output through PnP-RANSAC (“SCR”) and the fixed‑lag smoother that fuses IMU and SCR (“IMU + SCR”). A run is deemed failed when the smoother diverges. Our perception‑aware planner and Forward showed 100% of success rate, whereas the Constant showed 75% of success rate. Even on successful runs, our method consistently achieved lower translation and rotation errors. Our method lowers translation RMSE by 54% relative to the Constant and 15% relative to Forward; rotation RMSE is reduced by 40% and 31%, respectively. Similarly, it cuts the mean translation error by 45% versus Constant and 9% versus Forward, while the mean rotation error decreases by 35% and 27% compared with the same baselines.
Fig. 6(a) compares the estimated position trajectories and camera orientations produced by the two baselines and our planner for the “winter” reference trajectory. Fig. 6(b) plots the resulting translation and rotation errors over time. Our method tracks the path most accurately and achieves the lowest error, whereas Constant and Forward attains lower accuracy compared to our method.
@@
Consequently, the simulation results show that our perception-aware trajectory increases the localization accuracy.
@@
Computation time. We break down the runtime of the three main modules in our pipeline. The average runtimes for (1) SCR & PnP-RANSAC, (2) the fixed-lag smoothing, and (3) the trajectory optimizations are 62.2ms, 0.8ms, and 28.3ms, respectively. Given the execution periods of the optimization and SCR modules, the overall latency is sufficiently low for real-time implementation.
TABLE III: Ablation study. Translation and rotation RMSE and mean error are given in cm/, respectively.
IMU+SCR SCR Method RMSE $\downarrow$ Mean $\downarrow$ Median $\downarrow$ Mean $\downarrow$ Vanilla 22.6/4.4 20.3/3.9 10.1/0.9 13.7/1.2  + OW 24.0/3.8 21.5/3.1 10.2/0.8 14.1/1.2  + OW + OR 21.5/2.7 19.2/2.4 9.8/1.1 13.4/1.2  + OW + OR + SC 22.8/3.6 20.6/3.1 10.1/1.1 13.7/1.2
Ablation study. To assess the impact of uncertainty modeling, we progressively activated three components and reported the results in Table III. Vanilla uses the perception‑aware planner with the weight in Eq. (10) set to 1.0 and applies no entropy filtering in either the optimization stage or SCR. OW recovers the entropy weighting in Eq. (10). OR additionally discards high‑entropy pixels when sampling the $n_{f}$ scene coordinates for the trajectory optimization; this is our full method. SC further excludes high‑entropy points when sampling scene coordinates for PnP-RANSAC.
Enabling OW only reduces both rotation errors but increases the translation errors except the mean translation error of SCR. Introducing OR raises the median rotation errors of SCR, yet all other error metrics reach their lowest values, showing that the combination of entropy-based cost weights and discarding high‑entropy samples during optimization yields the best localization accuracy. Adding SC causes a performance drop, suggesting that filtering high-entropy pixels may not be beneficial for PnP-RANSAC. Therefore, we demonstrate that utilizing entropy for the trajectory optimization increases the localization accuracy.
(a) Phyiscal environment
@@
Figure 7: Hardware-in-the-loop experiment. (7(a)) Drone and spaces for the hardware-in-the-loop experiment. (7(b)) Time histories of translation error (up) and rotation error (down) for Forward (red) and Ours (blue)

### v2

@@ IV-B Trajectory Optimization
We receive the current UAV pose as estimated by the fixed-lag smoother alongside reference waypoints and scene coordinates with their corresponding uncertainties. Given these inputs, position and yaw trajectories are optimized to track waypoints and orient reliable scene coordinates. In this section, we detail the perception‑aware trajectory optimization.
@@
where the knot vector is $\{t_{0},t_{1},\dots,t_{n+k+1}\}$, the control points are $\mathbf{q}_{i}=\{q_{x,i},q_{y,i},q_{z,i},q_{\psi,i}\}\in\mathbb{R}^{4}$, and $N_{i,k}(t)$ denote the B‑spline basis functions. Each knot span has equal duration $\Delta t=t_{i+1}-t_{i}$. For clamping, the first and last $k$ knots are repeated: $t_{0}=\dots=t_{k-1}$ and $t_{n+1}=\dots=t_{n+k+1}$. We fix the first control point to the current UAV pose and the last control point to the final waypoint within the planning horizon $T_{\mathrm{plan}}$. The other control points become decision variables during the optimization.
Optimization Cost. Similar to Bartolomei et al. [4], the total cost function, $C_{\mathrm{total}}$, is represented as the weighted sum of the cost functions for different purposes. To reduce computational load, we optimize the position and yaw separately by minimizing two disjoint sub-objectives rather than the full $C_{total}$:
$\begin{gathered}C_{\mathrm{total}}^{\mathbf{p}}=\lambda_{\mathrm{wp}}C_{wp}^{\mathbf{p}}+\lambda_{\mathrm{eq}}C_{\mathrm{eq}}^{\mathbf{p}}+\lambda_{\mathrm{ie}}C_{\mathrm{ie}}^{\mathbf{p}}+\lambda_{s}C_{s}^{\mathbf{p}}\\ C_{\mathrm{total}}^{\psi}=\lambda_{\mathrm{fov}}C_{\mathrm{fov}}^{\psi}+\lambda_{\mathrm{eq}}C_{\mathrm{eq}}^{\psi}+\lambda_{\mathrm{ie}}C_{\mathrm{ie}}^{\psi}+\lambda_{s}C_{s}^{\psi}.\end{gathered}$ (5)
where the superscripts, $\mathbf{p}$ and $\psi$, indicate that the terms are evaluated on the position and yaw trajectory, respectively. For brevity, we will drop the superscripts in the following. The coefficients $\lambda_{(\cdot)}$ balance five components: $C_{\mathrm{wp}}$ penalizes deviation from the given waypoints, $C_{\mathrm{fov}}$ rewards keeping reliable scene coordinates within the camera’s field-of-view (FOV), $C_{\mathrm{eq}}$ enforces the trajectory starts from the given initial velocity, acceleration, and jerk, $C_{\mathrm{ie}}$ enforces the specified velocity and acceleration bounds, and $C_{s}$ promotes smoothness of the trajectory.
Let $\{\mathbf{w}_{0},\dots,\mathbf{w}_{r-1}\}$ denote all waypoints and let $\{s_{1},\dots,s_{r-2}\}$ be the prescribed arrival times for all but the first and last. The waypoint cost $C_{\mathrm{wp}}$ penalizes the squared Euclidean error between the position trajectory and each desired waypoint:
$C_{\mathrm{wp}}=\sum_{i=1}^{r-2}\,\sum_{o\in\{x,y,z\}}\bigl(q_{o}(s_{i})-w_{o,i}\bigr)^{2}.$ (6)
Our formulation of the FOV cost $C_{\mathrm{fov}}$ is inspired by Murali et al. [30]. Whereas the original cost encourages the camera to track as many covisible features as possible within its frustum, our version instead rewards keeping low‑uncertainty scene coordinates in view regardless of covisibility.
@@
Next, we need to build a differentiable indicator function describing whether scene coordinates $\mathbf{v}^{C}_{i,j}$ are within the camera frustum. For this purpose, we consider the displacement vectors from the optical center to each of the four image‑plane corners:
$\begin{split}\mathbf{z}_{TR}&=\big[(w-c_{x})l_{x}\quad(-c_{y})l_{y}\quad f\big]^{\top}\\ \mathbf{z}_{LR}&=\big[(w-c_{x})l_{x}\quad(h-c_{y})l_{y}\quad f\big]^{\top}\\ \mathbf{z}_{TL}&=\big[(-c_{x})l_{x}\quad(-c_{y})l_{y}\quad f\big]^{\top}\\ \mathbf{z}_{LL}&=\big[(-c_{x})l_{x}\quad(h-c_{y})l_{y}\quad f\big]^{\top}.\end{split}$ (7)
Here, $w$ and $h$ are the image width and height in pixels, $(c_{x},c_{y})$ is the principal point in pixels, $l_{x}$ and $l_{y}$ are the physical pixel sizes $(m/\text{pix})$, and $f$ is the focal length in meters. The normal vectors of the four side half planes of the camera frustum are $\mathbf{z}_{\mathrm{TR}}\!\times\!\mathbf{z}_{\mathrm{LR}}$, $\mathbf{z}_{\mathrm{TL}}\!\times\!\mathbf{z}_{\mathrm{TR}}$, $\mathbf{z}_{\mathrm{LL}}\!\times\!\mathbf{z}_{\mathrm{TL}}$, and $\mathbf{z}_{\mathrm{LR}}\!\times\!\mathbf{z}_{\mathrm{LL}}$. Using these normals, we construct a soft indicator vector
$\mathbf{o}_{i,j}=\frac{1}{2}\begin{bmatrix}1+\tanh\!\left((\mathbf{z}_{\mathrm{TR}}\!\times\!\mathbf{z}_{\mathrm{LR}})\!\cdot\!\mathbf{v}^{C}_{i,j}/s\right)\\[2.0pt] 1+\tanh\!\left((\mathbf{z}_{\mathrm{TL}}\!\times\!\mathbf{z}_{\mathrm{TR}})\!\cdot\!\mathbf{v}^{C}_{i,j}/s\right)\\[2.0pt] 1+\tanh\!\left((\mathbf{z}_{\mathrm{LL}}\!\times\!\mathbf{z}_{\mathrm{TL}})\!\cdot\!\mathbf{v}^{C}_{i,j}/s\right)\\[2.0pt] 1+\tanh\!\left((\mathbf{z}_{\mathrm{LR}}\!\times\!\mathbf{z}_{\mathrm{LL}})\!\cdot\!\mathbf{v}^{C}_{i,j}/s\right)\\[2.0pt] 1+\tanh\!\left((\mathbf{v}^{C}_{i,j}-f\mathbf{e}_{3})\!\cdot\!\mathbf{e}_{3}/s\right)\end{bmatrix}$, (8)
where $\mathbf{e}_{3}=[0,0,1]^{\top}$ and $s>0$ are the camera’s forward axis and a smoothing constant, respectively. For the $j$-th scene coordinate at time $t=s_{i}$, the desired scalar indicator is defined as the product of the elements of $\mathbf{o}_{i,j}$:
@@
$C_{fov}=-\sum_{i=1}^{r-1}\sum_{j=1}^{n_{f}}\exp(-a_{\mathrm{fov}}H_{\theta,j})F(\mathbf{c}_{i},\mathbf{v}_{j}).$ (10)
Here, $n_{f}$ denotes the number of selected scene coordinates. These coordinates are randomly chosen among those with entropy values below a predefined threshold. The weight $\exp(-a_{\mathrm{fov}}H_{\theta,j})$ depends on the entropy $H_{\theta,j}$ from Eq. (3); $a_{\mathrm{fov}}>0$ is constant. Lower entropy, i.e., higher certainty, yields a larger weight, encouraging the yaw trajectory to keep such scene coordinates in view.
@@
where $v_{o}(s_{i})$ and $a_{o}(s_{i})$ denote, respectively, the velocity and acceleration of coordinate $o$ at sample time $s_{i}$. The function $M_{u}(u_{o})=\max\big((u_{o}^{2}-u_{\max}^{2})^{2},0\big)$ is applied to both velocity $M_{v}(v_{o})$ and acceleration $M_{a}(a_{o})$.
@@
TABLE I: Localization accuracy for each translation in cm / rotation in . Columns show results for the four yaw strategies, Forward, Vanilla, FIM, and our proposed method (Ours). Accuracy is reported for raw SCR output (“SCR”) and after fusion with IMU data through the fixed-lag smoother (“IMU+SCR”). Best results are highlighted in bold, and second-best results are highlighted in underline.
@@ V Experiments
We begin by evaluating our approach in simulation. Our evaluation is based on two key assumptions: (A1) lower uncertainty implies higher scene‑coordinate accuracy, which we validate empirically in Section V-B, and (A2) uncertainty varies smoothly over the trajectory since it is spatially correlated, allowing its use within short planning horizons, as supported by Fig. 1 and the supplementary video. Using only current scene coordinates for short horizons rather than predicting future features is justified by the spatial correlation of the uncertainty.
Our two goals are: (a) to verify that the entropy defined in Eq. (3) provides an appropriate measure of uncertainty, and (b) to demonstrate the effectiveness of our perception‑aware trajectory planner. First, we empirically examine how the proposed entropy compares with conventional aleatoric and epistemic uncertainty measures. We then benchmark our planner against baselines. We also conduct hardware-in-the-loop experiments to validate system-level integration with onboard IMU data and simulated visual input.
Figure 4: Plots showing normalized uncertainty and scene coordinate error of E-SCRNet predictions. (left) Mean and standard deviation of L2 error (right) Density histogram.
Figure 5: Qualitative results of E-SCRNet for a test image. (upper left) Aleatoric Uncertainty (upper middle) Epistemic Uncertainty (upper right) Entropy (lower left) L2 Error of predicted scene coordinates (lower middle) Predicted scene coordinates (lower right) Input image.
(a) winter
(b) ampersand
Figure 6: Comparison of baseline planners (Forward, Vanilla, FIM) with the proposed method on “winter” and “ampersand” trajectories. (left) Time histories of translation and rotation errors of the fixed-lag smoother for Forward (blue), Vanilla (orange), FIM (green), and Ours (red) over the full duration of each trajectory. (right) Top‑down view showing reference trajectory (black dotted), the predicted trajectories (colored), camera orientations (green arrows) for one repetition of each trajectory.
@@ V-A Setup
We run photorealistic simulations using the FlightGoggles simulator on an Intel Core i9-7920X CPU [15]. The simulated camera resolution is fixed at $480\times 480$ pixels, with an FoV of $90^{\circ}$. We train the E‑SCRNet following the architecture in [7] for each of the two environments stata1 and stata2 as shown in Fig. 1.
To highlight the advantages brought by our perception‑aware planner, we deliberately use a small training set captured along a single trajectory (“picasso”) from the Blackbirds dataset [3], trained for only 100k iterations. We set the regularization coefficient $\rho=10^{-2}$. Following [7], we train the network using Adam with a learning rate of $10^{-4}$ and a batch size of 1, completing in 14 minutes on one NVIDIA RTX 6000 Ada GPU.
[…]
For the hardware-in-the-loop experiments, the quadrotor flies within an arena while onboard IMU data is logged and the camera stream is simulated with images rendered from the virtual stata1 scene, as shown in Fig. 7(a). The quadrotor’s pose is tracked using a motion capture system. An external machine equipped with an NVIDIA RTX A5000 GPU processes simulator images and executes SCR inference, smoothing, and optimization, transmitting control commands wirelessly to the UAV.
@@ V-B Uncertainty Representation
We collect scene coordinate predictions and their uncertainty maps from E‑SCRNet while the UAV follows a circular trajectory. Fig. 4 analyzes the relationship between the L2 prediction error and three uncertainty estimates: aleatoric, epistemic, and entropy. We clip the largest 3% of uncertainty values and normalize the remainder to $[0,1]$ for visual clarity.
As shown in Fig. 4, the mean error increases with increasing entropy, whereas its correlation with the others is weaker. Moreover, the error’s standard deviation declines as entropy decreases, while it remains high across the full range of the other measures. This empirical evidence directly validates assumption A1 stated at the beginning of this section. Fig. 4 further reveals that entropy is less skewed toward zero than the other uncertainties, indicating that it provides a more informative and well‑spread metric. Fig. 5 shows qualitative uncertainty and error maps for a representative test image.
@@
Comparison to baselines. Previous VIO-based perception-aware planners [30, 4] use an FOV cost which directs the camera toward visual features tracked over continuous frames in order to improve VIO accuracy. While this is similar to our proposed method, these approaches crucially do not make use of any uncertainty metric. In contrast, the proposed method leverages the uncertainties from our E-SCRNet to inform selection of reliable scene coordinates. To demonstrate the effectiveness of this additional uncertainty information, we compare our planner with three yaw strategies: First, Forward, which keeps the camera aligned with the velocity vector. Second, Vanilla, which employs the same perception-aware trajectory optimization as our proposed method but without incorporating uncertainty from E-SCRNet. Specifically, Vanilla uses a fixed weight in Eq. (10) of 1.0 and does not filter high-entropy pixels during yaw optimization. This baseline provides behavior equivalent to that from previous strategies [30, 4], while also making use of SCR for fairness in the comparison. Finally, Fisher information matrix (FIM) employs the same perception-aware trajectory optimization but replaces the entropy weight in Eq. (10) with the trace of the FIM introduced in Zhang et al. [40]. We evaluate four trajectories from the Blackbird dataset per site, “clover”, “patrick”, “thrice”, and “winter” in stata1, and “ampersand”, “dice”, “halfmoon”, and “sid” in stata2. Each trajectory is flown multiple times.
Table I reports localization accuracy for both the SCR output through PnP-RANSAC (“SCR”) and the fixed‑lag smoother that fuses IMU and SCR (“IMU+SCR”). Our method consistently outperforms the baselines on “IMU+SCR” localization, reducing translation RMSE by 15.0%, 4.9%, and 23.5% and rotation RMSE by 30.8%, 38.6%, and 41.3% compared to Forward, Vanilla, and FIM, respectively. Mean translation errors are reduced by 9.4%, 5.4%, and 19.7%, while mean rotation errors decrease by 27.3%, 38.5%, and 35.1% relative to the baselines. For “SCR” output, our method achieves the lowest mean errors compared to all baselines. Fig. 6 visualizes the estimated positions and camera orientations for selected trajectories, along with the histogram of translation and rotation, providing a visual representation of the quantitative results presented in Table I.
@@
Consequently, the simulation results validate that our perception-aware trajectory planner improves localization accuracy compared to baseline methods.
@@
Computation time. We break down the runtime of the main modules in our pipeline. The average runtimes for (1) SCR & PnP-RANSAC, (2) the fixed-lag smoothing, and (3) the trajectory optimizations are 62.2ms, 0.8ms, and 28.3ms, respectively. Given the execution periods of the optimization and SCR modules, the overall latency is sufficiently low for real-time implementation.
(a) Physical environment
@@
Figure 7: Hardware-in-the-loop experiment. (7(a)) Drone and spaces for the hardware-in-the-loop experiment. (7(b)) Time histories of translation error (up) and rotation error (down) for Forward and our method.